OpenStudy (anonymous):

y = x + 4 y = –2x + 1 I need this graphed

OpenStudy (anonymous):

Ok so you have a guess, with reasoning?

OpenStudy (texaschic101):

do you know how to find the x and y intercepts ?

OpenStudy (anonymous):

I just dont have a graphing calculator on me..

OpenStudy (texaschic101):

Can you graph it without the calculator ?

OpenStudy (anonymous):

And no, I forgot how to find the x and y intercepts

OpenStudy (texaschic101):

in y = mx + b form, the m is the slope and the b is the y intercept. so in the first one...y = x + 4....the slope is 1 and the y intercept is 4 or (0,4) to find the x intercept, sub in 0 for y y = x + 4 0 = x + 4 -4 = x....the x intercept is (-4,0) so plot your intercepts (0,4) and (-4,0) start at x intercept (0,-4) and go up 1, and to the right 1, then up 1, and to the right 1 and you should cross the y intercept at (0,4) Thats takes care of the first line Can you do the second line ?

OpenStudy (anonymous):

Probably

OpenStudy (texaschic101):

if you want to see the graph, go to mathway.com and it will show it

OpenStudy (texaschic101):

you probably already know this, but for the solution...it is where the 2 lines intersect

OpenStudy (anonymous):

I've been trying to do that for the past 20 minutes and its not working

OpenStudy (texaschic101):

whats not working ?

OpenStudy (anonymous):

It's not showing me the graph It will give me one line but not the other

OpenStudy (texaschic101):

Its giving me the other line....just type in the other line right below the first line you typed in ....solution is (-1,3)

OpenStudy (texaschic101):

but it did graph it

OpenStudy (texaschic101):

or you can go to Quickmath.com....thats another site where you can graph it